# On the complementation of spaces of $\mathcal{I}$-null sequences

### v1

@@ 5. When is $c_{0,\mathcal{I}}$ complemented in $\ell_{\infty}$?
Thus, $Q({\bf x})=\sum_{m\in F}(\mathcal{I}_{m}^{*}-\lim{\bf x})\chi_{V_{m}}$ when ${\bf x}=\sum_{i=1}^{k}c_{i}\chi_{A_{i}}$. The general case follows since the linear span of the set $\{\chi_{A}\,\colon\,A\subseteq\mathbb{N}\}$ is dense in $\ell_{\infty}$ (see [23, Proposition 2.5]). Therefore, $Q$ has the same form as given in (5.1). ∎
@@
We know that if $\mathcal{I}$ is an $\omega$-maximal ideal, then $\mathfrak{ad}(\mathcal{I}^{+})\leq\aleph_{0}$, as demonstrated earlier in the introduction to this section (see also [22]). The proof presented here shows also that if $\mathcal{I}$ is the intersection of a countable family of complemented ideals, then $\mathfrak{ad}(\mathcal{I}^{+})\leq\aleph_{0}$.
Now, we present examples of ideals that are not complemented. If $\mathcal{I}$ and $\mathcal{J}$ are ideals, their Fubini product ×$\mathcal{I}\times\mathcal{J}$ is the ideal on $\mathbb{N}\times\mathbb{N}$ defined by
@@
Regarding the Baire property of the Fubini product, in [11] was shown that ×$\mathcal{I}\times{\sf Fin}$ has the Baire property (hence, it is meager) for any ideal $\mathcal{I}$. On the other hand, they also showed that ${\sf Fin}\times\mathcal{I}$ has the Baire property exactly when $\mathcal{I}$ has it.
@@
    $\mathfrak{ad}({\sf Fin}\times\mathcal{I})=\mathfrak{ad}(\mathcal{I}\times{\sf Fin})=2^{\aleph_{0}}$. In particular, ${\sf Fin}\times\mathcal{I}$ and ×$\mathcal{I}\times{\sf Fin}$ are not complemented.
@@
To show that $\mathfrak{ad}(\mathcal{I}\times{\sf Fin})=2^{\aleph_{0}}$, we know that ×$\mathcal{I}\times{\sf Fin}$ has the Baire property by [11, Proposition 2.6]. Consequently, it is meager by Theorem 2.1. Therefore, we have $\mathfrak{ad}(\mathcal{I}\times{\sf Fin})=2^{\aleph_{0}}$ by Lemma 5.1. Now, from [18] it follows that ×$\mathcal{I}\times{\sf Fin}$ is not complemented.
(2) According to [23, Theorem 5.4], we have the isometry $c_{0,\mathcal{I}^{\omega\perp}}\cong c_{0}(c_{0,\mathcal{I}^{\perp}})$. By the main result of [5], the space $c_{0}(c_{0,\mathcal{I}^{\perp}})$ contains a complemented copy of $c_{0}$. Consequently, by Proposition 5.16, $c_{0,\mathcal{I}^{\omega\perp}}$ is not complemented in $\ell_{\infty}$. ∎
@@
The only examples of complemented ideals that we know are the strongly $\omega$-maximal ones. The only strict arrows we know are the ones mentioned in the diagram. To finish, we present the examples showing that the bottom arrows are strict. Example 3.8 shows that the first arrow is strict. For the second arrow, an example is given by [15, Lemma 2.2, Corollary 4.1]. Finally, for the last arrow, we mentioned above ${\sf Fin}\times\mathcal{I}$ is not meager whenever $\mathcal{I}$ is a maximal ideal, but $\mathfrak{ad}({\sf Fin}\times\mathcal{I})=2^{\aleph_{0}}$.
@@ Acknowledgments
We would like to thank Alan Dow for suggesting the Example 3.8. We also thank the partial support given by the VIE-grant 4248 of Universidad Industrial de Santander.

### v2

@@ 5. Structural properties of complemented ideals
Thus, $Q({\bf x})=\sum_{m\in F}(\mathcal{I}_{m}^{*}-\lim{\bf x})\chi_{V_{m}}$ when ${\bf x}=\sum_{i=1}^{k}c_{i}\chi_{A_{i}}$. The general case follows since the linear span of the set $\{\chi_{A}\,\colon\,A\subseteq\mathbb{N}\}$ is dense in $\ell_{\infty}$ (see [25, Proposition 2.5]). Therefore, $Q$ has the same form as given in (5.2). ∎
Now, in the same spirit of Theorems 5.2 and 5.11, we give a characterization of strongly $\omega$-maximal ideals in terms of the quotient $\ell_{\infty}/c_{0,\mathcal{I}}$.
Theorem 5.19.
Let $\mathcal{I}$ be an ideal on $\mathbb{N}$. The following statements are equivalent:
  1. (1)
    $K_{\mathcal{I}}$ is homeomorphic to $\beta\mathbb{N}$;
  2. (2)
    $\ell_{\infty}/c_{0,\mathcal{I}}$ is isometric to $\ell_{\infty}$;
  3. (3)
    $\ell_{\infty}/c_{0,\mathcal{I}}$ is Banach lattice isomorphic to $\ell_{\infty}$;
  4. (4)
    $\mathcal{I}$ is strongly $\omega$-maximal.
  5. (5)
    $c_{0,\mathcal{I}}$ is strongly complemented.
Proof.
$(1)\Longleftrightarrow(2)$ It is classical result that $C(\beta\mathbb{N})$ is isometric to $\ell_{\infty}$ and $C(K_{\mathcal{I}})$ is isometric to $\ell_{\infty}/c_{0,\mathcal{I}}$ by Theorem 5.1. Thus, the claim follows from the classical Banach-Stone theorem [1, Theorem 4.1.5].
$(1)\Longleftrightarrow(3)$ is shown analogously using a well known result of Kaplansky [17].
$(4)\Longleftrightarrow(5)$ is proved in Theorems 5.14 and 5.16.
$(4)\Rightarrow(3)$. Suppose that $\mathcal{I}$ is strongly $\omega$-maximal. By Proposition 3.1 and Lemma 2.3, there is a partition $\{A_{n}\,\colon\,n\in\mathbb{N}\}$ such that $A_{n}\in\mathcal{I}_{n}^{*}$ for every $n\in\mathbb{N}$ and $A\in\mathcal{I}$ if and only if $A\cap A_{n}\in\mathcal{I}_{n}$ for each $n\in\mathbb{N}$.
Claim 5.20.
The map
$\Psi\colon\ell_{\infty}$ $\to\ell_{\infty}((\ell_{\infty}/c_{0,\mathcal{I}_{m}})_{m\in\mathbb{N}})$
${\bf x}=(x_{n})$ $\mapsto((x_{n}\chi_{A_{m}}(n))+c_{0,\mathcal{I}_{m}})$
is an onto Banach lattice homomorphism and $\ker\Psi=c_{0,\mathcal{I}}$.
In fact, from the definition, it is clear that $\Psi$ is linear and preserves the lattice operations. To check continuity, we will use the following formula proved in [25, Lemma 5.7]. For ${\bf x}=(x_{n})\in\ell_{\infty}$ we have
$\|\Psi({\bf x})\|=\sup_{m\in\mathbb{N}}\|(x_{n}\chi_{A_{m}}(n))+c_{0,\mathcal{I}_{m}}\|=\sup_{m\in\mathbb{N}}\bigg(\mathcal{I}_{m}-\lim|x_{n}\chi_{A_{m}}(n)|\bigg)\leq\|{\bf x}\|.$
We will verify that $\Psi$ is onto. Let $({\bf y}^{m}+c_{0,\mathcal{I}_{m}})\in\ell_{\infty}((\ell_{\infty}/c_{0,\mathcal{I}_{m}})_{m\in\mathbb{N}})$ be given. For each $m\in\mathbb{N}$, let $a_{m}=\mathcal{I}_{m}-\lim{\bf y}^{m}$, which exists because $\mathcal{I}_{m}$ is maximal. Then, for every $m\in\mathbb{N}$,
${\bf y}^{m}+c_{0,\mathcal{I}_{m}}=a_{m}{\bf 1}+c_{0,\mathcal{I}_{m}}=a_{m}\chi_{A_{m}}+c_{0,\mathcal{I}_{m}}.$
Define ${\bf x}=(x_{n})\in\ell_{\infty}$ by $x_{n}=a_{m}$ whenever $n\in A_{m}$. Hence,
$\Psi({\bf x})=((x_{n}\chi_{A_{m}}(n))+c_{0,\mathcal{I}_{m}})=((a_{m}\chi_{A_{m}}(n))+c_{0,\mathcal{I}_{m}})=({\bf y}^{m}+c_{0,\mathcal{I}_{m}}).$
and therefore $\Psi$ is surjective.
Next, we prove that $\ker\Psi=c_{0,\mathcal{I}}$. Suppose that ${\bf x}=(x_{n})\in\ell_{\infty}$ satisfies that $\Psi({\bf x})={\bf 0}$. Then $((x_{n}\chi_{A_{m}}(n))\in c_{0,\mathcal{I}_{m}})$ for each $m\in\mathbb{N}$. Let $\varepsilon>0$ be given. By definition, $A(((x_{n}\chi_{A_{m}}(n)),\varepsilon)=A({\bf x},\varepsilon)\cap A_{m}\in\mathcal{I}_{m}$ for every $m\in\mathbb{N}$. By Lemma 2.3, it follows that $A({\bf x},\varepsilon)\in\mathcal{I}$. Since $\varepsilon>0$ was arbitrary, we conclude that ${\bf x}\in c_{0,\mathcal{I}}$.
Finally, the map $\Psi$ induces a Banach lattice isomorphism from $\ell_{\infty}/c_{0,\mathcal{I}}$ onto $\ell_{\infty}((\ell_{\infty}/c_{0,\mathcal{I}_{m}})_{m\in\mathbb{N}})$. Since $\dim\ell_{\infty}/c_{0,\mathcal{I}_{m}}=1$ for each $m\in\mathbb{N}$, we have that $\ell_{\infty}/c_{0,\mathcal{I}_{m}}$ is Banach lattice isomorphic to $\mathbb{R}$. By Remark 5.4, $\ell_{\infty}((\ell_{\infty}/c_{0,\mathcal{I}_{m}})_{m\in\mathbb{N}})$ is Banach lattice isomorphic to $\ell_{\infty}$. Consequently, $\ell_{\infty}/c_{0,\mathcal{I}}$ is Banach lattice isomorphic to $\ell_{\infty}$.
$(3)\Longrightarrow(4)$ Suppose that $\ell_{\infty}/c_{0,\mathcal{I}}$ is Banach lattice isomorphic to $\ell_{\infty}$ and let $T\colon\ell_{\infty}/c_{0,\mathcal{I}}\to\ell_{\infty}$ be a Banach lattice isomorphism. For each $n\in\mathbb{N}$, let ${\bf x}_{n}\in\ell_{\infty}$ be such that $T({\bf x}_{n}+c_{0,\mathcal{I}})=e_{n}$. Since ${\bf x}_{n}\not\in c_{0,\mathcal{I}}$, there are $\varepsilon_{n}>0$ and $A_{n}\not\in\mathcal{I}$ satisfying $\varepsilon_{n}\chi_{A_{n}}\leq|{\bf x}_{n}|$. Thus, $T(\varepsilon_{n}\chi_{A_{n}}+c_{0,\mathcal{I}})\leq T({\bf x}_{n}+c_{0,\mathcal{I}})=e_{n}$. Hence, $T(\chi_{A_{n}}+c_{0,\mathcal{I}})=a_{n}e_{n}$ for some $a_{n}>0$.
Claim 5.21.
If $m,n\in\mathbb{N}$ and $m\neq n$, then $A_{n}\cap A_{m}\in\mathcal{I}$.
Indeed, we have $\chi_{A_{m}\cap A_{n}}+c_{0,\mathcal{I}}=(\chi_{A_{m}}+c_{0,\mathcal{I}})\wedge(\chi_{A_{n}}+c_{0,\mathcal{I}})$. Hence,
$T(\chi_{A_{m}\cap A_{n}}+c_{0,\mathcal{I}})=T(\chi_{A_{m}}+c_{0,\mathcal{I}})\wedge T(\chi_{A_{n}}+c_{0,\mathcal{I}})=a_{m}e_{m}\wedge a_{n}e_{n}={\bf 0}.$
Claim 5.22.
For each $n\in\mathbb{N}$, $\mathcal{I}\restriction A_{n}$ is maximal.
Fix $n\in\mathbb{N}$. Let $B\subseteq A_{n}$ be such that $B\not\in\mathcal{I}$. Then, $T(\chi_{B}+c_{0,\mathcal{I}})\leq T(\chi_{A_{n}}+c_{0,\mathcal{I}})=a_{n}e_{n}$. Thus, $T(\chi_{B}+c_{0,\mathcal{I}})=b_{n}e_{n}$ for some $0<b_{n}\leq a_{n}$. Consequently, $T((\chi_{A_{n}}-\frac{a_{n}}{b_{n}}\chi_{B})+c_{0,\mathcal{I}})={\bf 0}$. It follows that $A_{n}\setminus B\in\mathcal{I}$ because of $A_{n}\setminus B\subset A({\bf z},1/2)$ where ${\bf z}=\chi_{A_{n}}-\frac{a_{n}}{b_{n}}\chi_{B}\in c_{0,\mathcal{I}}$.
Claim 5.23.
$B\in\mathcal{I}$ if, and only if, $B\cap A_{n}\in\mathcal{I}$ for each $n\in\mathbb{N}$.
Suppose that $B\cap A_{n}\in\mathcal{I}$ for each $n\in\mathbb{N}$ and let $T(\chi_{B}+c_{0,\mathcal{I}})={\bf y}=(y_{n})$. Fix $m\in\mathbb{N}$. Notice that ${\bf y}\geq 0$, and
${\bf 0}=T(\chi_{B\cap A_{m}}+c_{0,\mathcal{I}})=T(\chi_{B}+c_{0,\mathcal{I}})\wedge T(\chi_{A_{m}}+c_{0,\mathcal{I}})={\bf y}\wedge(a_{m}e_{m})=\min\{y_{m},a_{m}\}e_{m}.$
Thus, $y_{m}=0$. Since $m\in\mathbb{N}$ was arbitrary, we conclude that ${\bf y}={\bf 0}$. So, $B\in\mathcal{I}$.
Now, following the argument used in the proof of Lemma 4.5, we may assume that $A_{m}\cap A_{n}=\emptyset$ for every $m,n\in\mathbb{N}$ with $m\neq n$. Define $\mathcal{I}_{n}=\mathcal{I}\restriction A_{n}\sqcup\mathcal{P}(A_{n}^{c})$ for each $n\in\mathbb{N}$. By (1) of Lemma 2.2 and Claim 5.22, $\mathcal{I}_{n}$ is maximal for all $n\in\mathbb{N}$. It follows from Claim 5.23 that $\mathcal{I}=\bigcap_{n\in\mathbb{N}}\mathcal{I}_{n}$. Therefore, by Proposition 3.1, $\mathcal{I}$ is strongly $\omega$-maximal. ∎
6. Relative complementation and AD-families
In this section, we study when $c_{0,\mathcal{I}}$ is complemented in $c_{0,\mathcal{J}}$ when $\mathcal{I}\subseteq\mathcal{J}$. Also, we discuss the existence of $\mathcal{I}$-AD families and their connection with the complementation of $\mathcal{I}$.
$\mathfrak{ad}_{\mathcal{I}}(\mathcal{A})=\max\{|\mathcal{B}|\,\colon\,\mbox{$\mathcal{B}\subseteq\mathcal{A}$ is an $\mathcal{I}$-AD family}\}.$
We begin by stating a lemma, which is a well-known consequence of Theorem 2.1 (for a proof, see, for instance, [20]).
Lemma 6.1.
On the other hand, $\mathfrak{ad}(\mathcal{\mathcal{I}}^{+})\leq\aleph_{0}$ for every $\omega$-maximal ideal $\mathcal{I}$, this was implicitly shown by Plewik [24] and, for the sake of completion, we include a direct proof. Let $\mathcal{I}=\bigcap_{n}\mathcal{I}_{n}$ be an $\omega$-maximal ideal and suppose $\mathcal{B}\subseteq\mathcal{I}^{+}$ is an uncountable $\mathcal{I}$-AD family. Then, there is $n$ such that $|\mathcal{B}\cap\mathcal{I}_{n}^{+}|\geq 2$. If $A,B\in\mathcal{B}\cap\mathcal{I}_{n}^{+}$ are two different sets, then $A\cap B\in\mathcal{I}\subseteq\mathcal{I}_{n}$, which contradicts that $\mathcal{I}_{n}$ is maximal. Additionally, any $\omega$-maximal ideal is not measurable with respect to the usual product measure on $2^{\mathbb{N}}$ (see [16, section 4.1]).
The proof presented above shows that:
Corollary 6.2.
(Kania [15, Theorem A]) Let $\mathcal{I}$ be an ideal on $\mathbb{N}$. If $\mathcal{I}$ is complemented, then $\mathfrak{ad}(\mathcal{I}^{+})\leq\aleph_{0}$.
The following theorem extends Kania’s result to ideals $\mathcal{I}$ and $\mathcal{J}$ on $\mathbb{N}$ such that $\mathcal{I}\subsetneq\mathcal{J}$ and $c_{0,\mathcal{I}}$ complemented in $c_{0,\mathcal{J}}$.
Theorem 6.3.
Let $\mathcal{I}\subsetneq\mathcal{J}$ be proper ideals on $\mathbb{N}$. Suppose that $c_{0,\mathcal{I}}$ is kernel of an operator from $c_{0,\mathcal{J}}$ to $c_{0,\mathcal{J}}$. Then, $\mathfrak{ad}_{\mathcal{I}}(\mathcal{J}\cap\mathcal{I}^{+})\leq\aleph_{0}$. In particular, if $c_{0,\mathcal{I}}$ is complemented in $c_{0,\mathcal{J}}$, then $\mathfrak{ad}_{\mathcal{I}}(\mathcal{J}\cap\mathcal{I}^{+})\leq\aleph_{0}$.
Proof.
Let $Q\colon c_{0,\mathcal{J}}\to c_{0,\mathcal{J}}$ be a bounded linear operator such that $\ker Q=c_{0,\mathcal{I}}$. For $k,n\in\mathbb{N}$, let $\mathcal{A}_{n,k}$ be as in Lemma 5.10. Then, we have $\mathcal{J}\cap\mathcal{I}^{+}=\bigcup_{n,k}\mathcal{A}_{n,k}$ and $\mathfrak{ad}_{\mathcal{I}}(\mathcal{A}_{n,k})<\aleph_{0}$ for each $n,k$. Suppose $\mathcal{B}\subseteq\mathcal{J}\cap\mathcal{I}^{+}$ is an $\mathcal{I}$-AD family. Then, for every $n,k\in\mathbb{N}$, $\mathcal{B}\cap\mathcal{A}_{n,k}$ is finite. Thus, $\mathcal{B}$ is countable, and consequently $\mathfrak{ad}_{\mathcal{I}}(\mathcal{J}\cap\mathcal{I}^{+})\leq\aleph_{0}$. ∎
The next observation follows immediately from Lemma 6.1.
Corollary 6.4.
Let $\mathcal{I}\subsetneq\mathcal{J}$ be proper ideals on $\mathbb{N}$ such that $\mathcal{I}\restriction A$ is meager as a subset of $2^{A}$ for some infinite set $A\in\mathcal{J}$. Then, $\mathfrak{ad}_{\mathcal{I}}(\mathcal{J}\cap\mathcal{I}^{+})>\aleph_{0}$.
From Theorem 6.3 we get the following corollary. A similar result appears in [20, Corollary 1.5] about the space $c$ of convergent sequences.
Corollary 6.5.
Proposition 6.6.
Proof.
@@ 6. Relative complementation and AD-families
Proposition 6.8.
Proof.
Now, we present examples of ideals that are not complemented. If $\mathcal{I}$ and $\mathcal{J}$ are ideals, their Fubini product $\mathcal{I}\times\mathcal{J}$ is the ideal on $\mathbb{N}\times\mathbb{N}$ defined by
@@
Regarding the Baire property of the Fubini product, in [11] was shown that $\mathcal{I}\times{\sf Fin}$ has the Baire property (hence, it is meager) for any ideal $\mathcal{I}$. On the other hand, they also showed that $\sf Fin$$\times\mathcal{I}$ has the Baire property exactly when $\mathcal{I}$ has it.
@@
    $\mathfrak{ad}({\sf Fin}\times\mathcal{I})=\mathfrak{ad}(\mathcal{I}\times{\sf Fin})=2^{\aleph_{0}}$. In particular, $\sf Fin$$\times\mathcal{I}$ and $\mathcal{I}\times{\sf Fin}$ are not complemented.
@@
To show that $\mathfrak{ad}(\mathcal{I}\times{\sf Fin})=2^{\aleph_{0}}$, we know that $\mathcal{I}\times{\sf Fin}$ has the Baire property by [11, Proposition 2.6]. Consequently, it is meager by Theorem 2.1. Therefore, we have $\mathfrak{ad}(\mathcal{I}\times{\sf Fin})=2^{\aleph_{0}}$ by Lemma 6.1. Now, from [20] it follows that $\mathcal{I}\times{\sf Fin}$ is not complemented.
(2) According to [25, Theorem 5.4], we have the isometry $c_{0,\mathcal{I}^{\omega\perp}}\cong c_{0}(c_{0,\mathcal{I}^{\perp}})$. By the main result of [5], the space $c_{0}(c_{0,\mathcal{I}^{\perp}})$ contains a complemented copy of $c_{0}$. Consequently, by Proposition 6.8, $c_{0,\mathcal{I}^{\omega\perp}}$ is not complemented in $\ell_{\infty}$. ∎
@@
To conclude, we provide examples showing that all the implications in the diagram are strict. Example 3.8 shows that the first implication in the lower row is strict; consequently, the corresponding implication in the upper row is also strict. Moreover, there exists a complemented ideal that is not $\omega$-maximal [13, Example 2], which shows that the second implication in the upper row is strict. For the second implication in the lower row, an example is given by [16, Lemma 2.2 and Corollary 4.1]. A counterexample to the last downward implication between the two rows of the diagram is given in [13, Example 1]. Finally, for the last implication in the lower row, recall that $\sf Fin$$\times\mathcal{I}$ is not meager whenever $\mathcal{I}$ is a maximal ideal, but $\mathfrak{ad}({\sf Fin}\times\mathcal{I})=2^{\aleph_{0}}$.
@@ Acknowledgments
We would like to thank Alan Dow for suggesting the Example 3.8. We also thank for the partial support given by the VIE-grant 4248 of Universidad Industrial de Santander.